# Quantitative evaluation of methods to analyze motion changes in single-particle experiments

### v2

@@ Abstract
The analysis of live-cell single-molecule imaging experiments can reveal valuable information about the heterogeneity of transport processes and interactions between cell components. These characteristics are seen as motion changes in the particle trajectories. Despite the existence of multiple approaches to carry out this type of analysis, no objective assessment of these methods has been performed so far. Here, we have designed a competition to characterize and rank the performance of these methods when analyzing the dynamic behavior of single molecules. To run this competition, we have implemented a software library to simulate realistic data corresponding to widespread diffusion and interaction models, both in the form of trajectories and videos obtained in typical experimental conditions. The competition will constitute the first assessment of these methods, provide insights into the current limits of the field, foster the development of new approaches, and guide researchers to identify optimal tools for analyzing their experiments.
@@
random walk; heterogeneous diffusion; anomalous diffusion; trajectory analysis; time series; single-particle tracking; stochastic systems; deep learning; changepoint analysis
Introduction
Figure 1: Rationale for the challenge organization. a, The interactions of biomolecules in complex environments, such as the cell membrane, regulate physiological processes in living systems. These interactions produce changes in molecular motion that can be used as a proxy to measure interaction parameters. b-c, Time-lapse single-molecule imaging allows us to visualize these processes with high spatiotemporal resolution (b) and, in combination with single-particle tracking methods, provide trajectories of individual molecules (c). d-e, Analytical methods can be applied to imaging data, either raw (b) or processed in the form of trajectories (c), to infer interaction kinetics and quantify their dynamic properties at the ensemble (e.g., probability distributions, d) or single-trajectory level (e.g., changepoints, e).
@@ Introduction
Physiological processes occurring in living cells rely on encounters and interactions between molecules. Archetypal examples include gene regulation, transduction of biological signals, and protein delivery to specific locations. All these processes involve the active or passive transport of biomolecules in highly complex, time-varying, and far-from-equilibrium environments, such as the cell membrane (1(a)). One of the most powerful tools to study these transport phenomena is the combination of live-cell single-molecule imaging with single-particle tracking [1, 2] because it can provide the time and location when and where single events take place (1(b-c)). Alternative ensemble methods (e.g., fluorescence correlation spectroscopy or fluorescence recovery after photobleaching [3]) usually provide limited information because they lose track of crucial information when averaging out spatial and temporal fluctuations.
@@
Methods for single-molecule imaging and single-particle tracking have seen tremendous progress in the last decade, in terms of both experimental acquisition and data analysis [1, 2, 4, 5]. The abundance of experimental single-particle trajectories, encompassing molecules, protein complexes, vesicles, and organelles, has led to the development of numerous methods dedicated to the reliable detection of changes in their motion patterns (as summarized in Table 1). These changes serve as valuable indicators for the occurrence of interactions within the system. For instance, diffusing particles may exhibit variations in diffusion coefficients (due to processes like dimerization, ligand binding, or conformational changes) or shifts in their mode of motion (attributed to transient immobilization or confinement at specific scaffolding sites) (1(a)) [6]. These interactions can also result in deviations from standard Brownian motion, as characterized by Einstein’s free diffusion model, which includes a linear mean-squared displacement (MSD) and a Gaussian distribution of displacements [7]. This is the case, e.g., of spatiotemporal heterogeneities producing transient subdiffusion at specific timescales [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Other mechanisms can instead produce asymptotically-anomalous diffusion [20, 21, 22, 2]. Anomalous diffusion compatible with models such as fractional Brownian motion [23, 24, 25, 26, 27, 28], continuous-time random walk [29, 30], scaled Brownian motion [31], and Lévy walk [32] has been observed for telomers, macromolecular complexes, proteins, and organelles in living cells. Several approaches have been recently proposed to detect and quantify these behaviors [33, 34], also involving machine-learning techniques [35, 36, 37, 38, 39, 40, 41].
@@
To gain insights into the performance of methods to detect anomalous diffusion from individual trajectories, in 2021 we successfully ran the $1^{\rm st}$ AnDi Challenge [42]. The discussion that developed between members of diverse research communities working on biology, microscopy, single-particle tracking, and anomalous diffusion (including experimentalists, theoreticians, data analysts, and computer scientists) emphasized the necessity for deeper insights into biologically relevant phenomena. First, there is a need to evaluate methods to determine the switch between different diffusive behaviors, as often observed in experiments. Second, it is necessary to assess the methods’ crosstalk in detecting inherent anomalous diffusion from nonlinearity in the MSD due to motion constraints or heterogeneity. Third, there is a need to determine whether the bottleneck of the analysis process is at the level of the analysis of the single trajectories or associated with their extraction from the experimental videos. These needs shaped the design of the $2^{\rm nd}$ AnDi Challenge, defining its scope with a focus on characterizing and ranking the performance of methods that analyze changes of dynamic behavior. While we have retained the name of the $1^{\rm st}$ AnDi Challenge to build upon its already-established community, we would like to emphasize that the main focus of this $2^{\rm nd}$ AnDi Challenge is on revealing heterogeneity rather than anomalous diffusion. In the simulated datasets, anomalous diffusion will either emerge from heterogeneity itself or be intentionally introduced for evaluation purposes.
@@
There is a multitude of methods that have been designed to identify and characterize heterogeneous diffusion (Table 1). They can be classified based on the heterogeneity they aim to identify or on the kind of analysis they perform. We consider three heterogeneity classes that these methods aim to identify: (i) changes in the value of the diffusion coefficient $D$; (ii) changes in the anomalous diffusion exponent $\alpha$ (often classified as subdiffusion, diffusion, or superdiffusion); and (iii) changes in the phenomenological behavior associated with interactions with the environment (often classified as immobilization, confinement, (free) diffusion, and directed motion). While changes in the diffusion coefficient and in the phenomenological behavior have been widely reported, the exploration of changes in the anomalous diffusion exponent is a more recent development [43, 44, 45, 46], which is attracting increasing interest also from the theoretical point of view [47, 48, 49, 50]. The introduction of new methods for data analysis, as promoted by the Challenge, could push the performance for detecting subtle changes in these diffusion properties in systems where they have so far been overlooked. Along this line, it must be pointed out that the traditional analysis based on the calculation of the scaling exponent of the mean-squared displacement (MSD) can create some ambiguity between the last two classes. Just to provide an example, a particle performing Brownian diffusion in a confined region has an exponent $\alpha=1$ in terms of the generating motion, but its MSD features a horizontal asymptote at long times, corresponding to $\alpha=0$. In the following, we will refer to the exponent $\alpha$ as the characteristic feature of the generating motion.
@@
From the analysis point of view, we can identify two classes of methods: (i) ensemble methods, meant to determine characteristic features out of an ensemble of trajectories (1(d)) and (ii) single-trajectory methods, meant to identify changepoint (CP) locations through trajectory segmentation (1(e)). While most available methods rely on the analysis of trajectories obtained from video processing [51], recent advances in computer vision have led to methods capable of directly extracting information from raw movies without requiring the explicit extraction of trajectories [52, 53]. Each method has its own set of advantages and disadvantages, and its performance may depend on the specific problem under consideration. However, there is currently no universally accepted gold standard for determining which method to use to address each specific problem.
@@
To cater to these more advanced needs, we decided to run an open competition as the $2^{\rm nd}$ Anomalous Diffusion (AnDi) Challenge. The rationale described above shaped the scope of the challenge, defining the choice of the datasets and the design of the tasks. To rely on objective ground truth, we will assess the methods’ performance on simulated datasets inspired by models of diffusion and interactions documented in biological systems. Datasets will describe particles undergoing fractional Brownian motion (FBM, [54]) with piecewise-constant parameters. FBM-type motion has been widely observed in biological systems by means of microrheology, a technique that uses large tracer particles as probes to study the properties of the environment [55]. Anomalous diffusion compatible with FBM has also been reported for telomers and macromolecular complexes in living cells [56, 23, 26, 24, 27, 25, 28, 20]. Beyond this evidence, in the context of the Challenge, FBM serves as a tool to enable the tuning of diffusion parameters. The combination of parameter values and interaction models might produce situations that do not correspond to previously documented biological scenarios but will be valuable to test the methods’ performance in a wide range of conditions. In biological experiments, other kinds of motion and even non-Gaussian behavior have been reported [21]. However, the choice of FBM does not limit the generality of the Challenge since other models of diffusion and non-Gaussian behavior can be obtained by properly tuning the parameters of the simulations. Datasets provided for the last phase of the competition will also include actual experiments for their comparative analysis with the Challenge methods (these data will not be used for the ranking).
@@
The standard and straightforward approach in live-cell single-molecule imaging primarily captures information related to lateral motion. In cases involving flat membranes or isotropic systems, employing 2D imaging and tracking techniques suffices for obtaining accurate motion-related parameters. However, when dealing with motion on non-flat surfaces or within anisotropic 3D environments, relying solely on 2D projections can result in critical information being overlooked, potentially leading to the misinterpretation of diffusion coefficients or the appearance of apparent anomalous diffusion effects [57, 58]. Consequently, drawing definitive conclusions under such circumstances should be avoided or approached with caution. To study motion occurring in 3D space, it is advisable to employ 3D tracking methods, such as off-focus imaging (i.e., the analysis of ring patterns in the defocused point spread function) [59], interference/holographic approaches [60], multifocus imaging [61], or point spread function engineering [62]. Although more challenging, these methods can measure also the motion along the axial dimension, facilitating a more accurate characterization. For the purposes of the Challenge, we have chosen to concentrate on studying changes in diffusion behavior occurring within a 2D context, driven by particle interactions of various types.
@@
While this challenge focuses on data from biological systems, the use of regime-switching detection and trajectory segmentation extends well beyond the domain of living cells. Particularly interesting applications also include, e.g., the analysis of biomedical signals [63], speech [64], traffic flows [65], seismic signals [66], econometrics [67, 68], ecology [69], and river flows [70].
Methods
Datasets and ground truth
@@ Methods
In order to benchmark the different methods on data with known ground truth, we rely on numerical simulations. We developed the andi-datasets Python package [71] to generate the required datasets to train and evaluate the various methods. Details about available functions can be found in the hosting repository [71].
@@
Particle motion is simulated according to fractional Brownian motion (FBM, [54]), a model that reproduces Brownian and anomalous diffusion processes by tuning the correlation of the increments through the Hurst exponent $H$. FBM is a Gaussian process with a covariance function
@@
where ${\rm E}[\cdot]$ denotes the expected value and $K$ is a constant with units ${\rm length^{2}\cdot time^{-2H}}$. In order to generalize FBM in two dimensions (2D), a trajectory ${\bf R}(t)$ is represented as ${\bf R}(t)=\{X(t),Y(t)\}$, where $X(t)$ and $Y(t)$ are independent FBM along the $x$ and $y$ axes, respectively [33]. The anomalous diffusion exponent is related to the Hurst exponent as $\alpha=2H$ [54], and the MSD for an unconstrained FBM in 2D scales with time $t$ as
@@
When $\alpha=1$, FBM reverts to Brownian motion and $K$ corresponds to the diffusion coefficient $D$. FBM can thus produce subdiffusion for $20<H<1/$ 2 ($10<\alpha<$ 1), Brownian diffusion for $H=1/2$ ($\alpha=1$), and superdiffusion for $11/2<H<$ 1 ($21<\alpha<$ 2).
We will consider the following physical models of motion and interactions(2(a)):
@@
[…]
A detailed description of the simulation procedure is presented in A. Table 2 summarizes all the parameter values used to generate the Pilot Dataset that can be downloaded at https://drive.google.com/drive/folders/1R41I5Y8ynXdXbUFLWOtQXx1nkUh3bWw3?usp=share_link. The Pilot Dataset provides typical numerical experiments that will be made available to the participants in the different phases of the challenge.
Datasets provided for the last phase of the competition will also include results from actual experiments that have reported the occurrence of heterogeneous diffusion but for which the ground truth is not established beyond any doubt. Therefore, these data will not contribute to the challenge scoring. Nevertheless, the predictions provided by different methods will be comparatively analyzed and discussed also with respect to the conclusions reached in the original publications. Together with the quantitative results obtained from simulations, these analyses will assess the applicability of the methods to real-world experimental data.
Competition design
To enable the assessment of the performance of previously established methods while fostering the development of new approaches and the participation from diverse disciplines, the challenge is organized along two tracks:
  • Track 1 — based on the analysis of raw videos.
  • Track 2 — based on the analysis of trajectories.
For each track, datasets are provided according to a hierarchical structure (2(b-c)) that includes:

### v3

@@ Abstract
The analysis of live-cell single-molecule imaging experiments can reveal valuable information about the heterogeneity of transport processes and interactions between cell components. These characteristics are seen as motion changes in the particle trajectories. Despite the existence of multiple approaches to carry out this type of analysis, no objective assessment of these methods has been performed so far. Here, we report the results of a competition to characterize and rank the performance of these methods when analyzing the dynamic behavior of single molecules. To run this competition, we implemented a software library that simulates realistic data corresponding to widespread diffusion and interaction models, both in the form of trajectories and videos obtained in typical experimental conditions. The competition constitutes the first assessment of these methods, providing insights into the current limitations of the field, fostering the development of new approaches, and guiding researchers to identify optimal tools for analyzing their experiments.
keywords:
@@
random walk, heterogeneous diffusion, anomalous diffusion, trajectory analysis, time series, single-particle tracking, stochastic systems, deep learning, changepoint analysis
1 Comments
This is the author’s version of the article published in Nature Communications under a Creative Commons Attribution 4.0 International License (CC BY 4.0). The final published version is available at https://doi.org/10.1038/s41467-025-61949-x.
2 Protocol registration
The Stage 1 protocol for this Registered Report was accepted-in-principle on 31st October 2023. The protocol, as accepted by the journal, can be found at https://doi.org/10.6084/m9.figshare.24771687.v1.
3 Introduction
@@
Physiological processes occurring in living cells rely on encounters and interactions between molecules. Archetypal examples include gene regulation, transduction of biological signals, and protein delivery to specific locations. All these processes involve the active or passive transport of biomolecules in highly complex, time-varying, and far-from-equilibrium environments, such as the cell membrane (Fig.˜1a). One of the most powerful tools to study these transport phenomena is the combination of live-cell single-molecule imaging with single-particle tracking [1, 2] because it can provide the time when and location where single events take place (Fig.˜1b,c). Alternative ensemble methods (e.g., fluorescence correlation spectroscopy or fluorescence recovery after photobleaching [3]) usually provide limited information because they lose track of crucial details when averaging out spatial and temporal fluctuations.
Figure 1: Rationale for the challenge organization. a, The interactions of biomolecules in complex environments, such as dimerization, ligand binding, or trapping at the cell membrane, regulate physiological processes in living systems. These interactions produce changes in molecular motion that can be used as a proxy to measure interaction parameters. b, c, Time-lapse single-molecule imaging allows us to visualize these processes with high spatiotemporal resolution (b) and, in combination with single-particle tracking methods, provide trajectories of individual molecules (c). d, e, Analytical methods can be applied to imaging data, either raw (b) or processed in the form of trajectories (c), to infer interaction kinetics and quantify their dynamic properties at the ensemble (e.g., probability distributions, d) or single-trajectory level (e.g., changepoints, e).
@@
Methods for single-molecule imaging and single-particle tracking have seen tremendous progress in the last decade, in terms of both experimental acquisition and data analysis [1, 2, 4, 5]. The abundance of experimental single-particle trajectories, encompassing molecules, protein complexes, vesicles, and organelles, has led to the development of numerous methods dedicated to the reliable detection of changes in their motion patterns (as summarized in Supp. Tab. 1). These changes serve as valuable indicators for the occurrence of interactions within the system. For instance, diffusing particles may exhibit variations in diffusion coefficients (due to processes like dimerization, ligand binding, or conformational changes) or shifts in their mode of motion (attributed to transient immobilization or confinement at specific scaffolding sites) (Fig.˜1a) [6]. These interactions can also result in deviations from standard Brownian motion, as characterized by Einstein’s free diffusion model, which includes a linear mean-squared displacement (MSD) and a Gaussian distribution of displacements [7]. This is the case, e.g., of spatiotemporal heterogeneities producing transient subdiffusion at specific timescales [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Other mechanisms can instead produce asymptotic anomalous diffusion [20, 21, 22, 2]. Anomalous diffusion compatible with models such as fractional Brownian motion [23, 24, 25, 26, 27, 28], continuous-time random walk [29, 30], scaled Brownian motion [31], and Lévy walk [32] has been observed for telomers, macromolecular complexes, proteins, and organelles in living cells. Several approaches have been recently proposed to detect and quantify these behaviors [33, 34], also involving machine-learning techniques [35, 36, 37, 38, 39, 40, 41].
@@
To gain insights into the performance of methods to detect anomalous diffusion from individual trajectories, in 2021 we successfully ran the $1^{\rm st}$ AnDi Challenge [42]. The discussion that developed between members of diverse research communities working on biology, microscopy, single-particle tracking, and anomalous diffusion (including experimentalists, theoreticians, data analysts, and computer scientists) emphasized the necessity for deeper insights into biologically relevant phenomena. First, it identified a need to evaluate methods to determine the switch between different diffusive behaviors, as often observed in experiments. Second, it highlighted the necessity to assess the methods’ crosstalk in detecting inherent anomalous diffusion from nonlinearity in the MSD due to motion constraints or heterogeneity. Third, it emphasized the importance of determining whether the bottleneck of the analysis process was at the level of the analysis of the single trajectory or associated with their extraction from experimental videos. These needs shaped the design of the $2^{\rm nd}$ AnDi Challenge, defining its scope with a focus on characterizing and ranking the performance of methods that analyze changes of dynamic behavior. While we retained the name of the $1^{\rm st}$ AnDi Challenge to build upon its already-established community, the $2^{\rm nd}$ AnDi Challenge focused mainly on revealing heterogeneity rather than anomalous diffusion. In the simulated datasets, anomalous diffusion emerged from heterogeneity itself or was intentionally introduced for evaluation purposes.
@@
A multitude of methods have been designed to identify and characterize heterogeneous diffusion (Supp. Tab. 1). They can be classified based on the heterogeneity they aim to identify or the kind of analysis they perform. We considered three heterogeneity classes that these methods aim to identify: (i) changes in the diffusion coefficient $D$; (ii) changes in the anomalous diffusion exponent $\alpha$ (often classified as subdiffusion, diffusion, or superdiffusion); and (iii) changes in the phenomenological behavior associated with interactions with the environment (often classified as immobilization, confinement, (free) diffusion, and directed motion). While changes in the diffusion coefficient and in the phenomenological behavior have been widely reported, the exploration of changes in the anomalous diffusion exponent is a more recent development [43, 44, 45, 46], which is attracting increasing interest also from the theoretical point of view [47, 48, 49, 50]. The introduction of new methods for data analysis, as promoted by the Challenge, had the objective to push the performance for detecting subtle changes in these diffusion properties in systems where they could have been overlooked. Along this line, it must be pointed out that the traditional analysis based on the calculation of the scaling exponent of the mean-squared displacement (MSD) can create some ambiguity between the last two classes. Just to provide an example, a particle performing Brownian diffusion in a confined region has an exponent $\alpha=1$ in terms of the generating motion, but its MSD features a horizontal asymptote at long times, corresponding to $\alpha=0$. In the following, we will refer to the exponent $\alpha$ as the characteristic feature of the generating motion.
@@
From the analysis point of view, we identified two classes of methods: (i) ensemble methods, meant to determine characteristic features out of an ensemble of trajectories (Fig.˜1d) and (ii) single-trajectory methods, meant to identify changepoint (CP) locations through trajectory segmentation (Fig.˜1e). While most available methods rely on the analysis of trajectories obtained from video processing [51], recent advances in computer vision have led to methods capable of directly extracting information from raw movies without requiring the explicit extraction of trajectories [52, 53]. Each method has its own set of advantages and disadvantages, and its performance may depend on the specific problem under consideration. However, there is no universally accepted gold standard for determining which method to use to address each specific problem.
@@
To cater to these more advanced needs, we ran an open competition as the $2^{\rm nd}$ Anomalous Diffusion (AnDi) Challenge. The rationale described above shaped the scope of the challenge, defining the choice of the datasets and the design of the tasks. To rely on an objective ground truth, we assessed the methods’ performance on simulated datasets inspired by models of diffusion and interactions documented in biological systems. These datasets describe particles undergoing fractional Brownian motion (FBM, [54]) with piecewise-constant parameters. FBM-type motion has been widely observed in biological systems by means of microrheology, a technique that uses large tracer particles as probes to study the properties of the environment [55]. Anomalous diffusion compatible with FBM has also been reported for telomers and macromolecular complexes in living cells [56, 23, 26, 24, 27, 25, 28, 20]. Beyond this evidence, in the context of the Challenge, FBM served as a tool to enable the tuning of diffusion parameters. The combination of parameter values and interaction models might produce situations that do not correspond to previously documented biological scenarios but will be valuable to test the methods’ performance in a wide range of conditions. In biological experiments, other kinds of motion and even non-Gaussian behavior have been reported [21]. However, the choice of FBM did not limit the generality of the Challenge since other models of diffusion and non-Gaussian behavior can be obtained by properly tuning the parameters of the simulations. Datasets provided for the last phase of the competition included motion with parameters inspired by actual experiments for their comparative analysis with the Challenge methods.
@@
The standard and straightforward approach in live-cell single-molecule imaging primarily captures information related to lateral motion. In cases involving flat membranes or isotropic systems, employing 2D imaging and tracking techniques suffices for obtaining accurate motion-related parameters. However, when dealing with motion on non-flat surfaces or within anisotropic 3D environments, relying solely on 2D projections can result in critical information being overlooked, potentially leading to the misinterpretation of diffusion coefficients or the appearance of apparent anomalous diffusion effects [57, 58]. Consequently, drawing definitive conclusions under such circumstances should be avoided or approached with caution. To study motion occurring in 3D space, it is advisable to employ 3D tracking methods, such as off-focus imaging (i.e., the analysis of ring patterns in the defocused point spread function) [59], interference/holographic approaches [60], multifocus imaging [61], or point spread function engineering [62]. Although more challenging, these methods can also measure the motion along the axial dimension, facilitating a more thorough characterization. For the purposes of the Challenge, we choose to concentrate on studying changes in diffusion behavior occurring within a 2D context, driven by particle interactions of various types.
@@
While this challenge focused on data inspired by biological systems, the use of regime-switching detection and trajectory segmentation extends well beyond the domain of living cells. Particularly interesting applications also include, e.g., the analysis of biomedical signals [63], speech [64], traffic flows [65], seismic signals [66], econometrics [67, 68], ecology [69], and river flows [70].
4 Results
4.1 Datasets and ground truth
@@
In order to benchmark the different methods on data with a known ground truth, we relied on numerical simulations. We developed the andi-datasets Python package [71] to generate the required datasets to train and evaluate the various methods. Details about available functions can be found in the hosting repository [71].
@@
Particle motion was simulated according to fractional Brownian motion (FBM, [54]), a model that reproduces Brownian and anomalous diffusion processes by tuning the correlation of the increments through the Hurst exponent $H$. FBM is a Gaussian process with a covariance function
@@
where $\rm E$$\cdot$] denotes the expected value and $K$ is a constant with units $\rm length^{2}\cdot time^{-2H}$}. In order to generalize FBM in two dimensions (2D), a trajectory ${\bf R}(t)$ is represented as ${\bf R}(t)=\{X(t),Y(t)\}$, where $X(t)$ and $Y(t)$ are independent FBM processes along the $x$ and $y$ axes, respectively [33]. The anomalous diffusion exponent is related to the Hurst exponent as $\alpha=2H$ [54], and the MSD for an unconstrained FBM in 2D scales with time $t$ as
@@
When $\alpha=1$, FBM reverts to Brownian motion and $K$ corresponds to the diffusion coefficient $D$. FBM describes subdiffusion for $0<H<1/2$ ($0<\alpha<1$), Brownian diffusion for $H=1/2$ ($\alpha=1$), and superdiffusion for $1/2<H<1$ ($1<\alpha<2$).
Figure 2: Physical models of interaction and structure of the simulated datasets. a, Examples of 2-dimensional trajectories undergoing interactions inducing changes in their motion. From left to right: single-state model (SSM) without changes of diffusion; multi-state model (MSM) with time-dependent changes between different diffusive states (red and blue); dimerization model (DIM) where a particle (red) selectively interacts with another particle (gray) and the two transiently co-diffuse with a different motion (blue trajectory); transient-confinement model (TCM) where a particle diffuses inside (blue) and outside (red) compartments with osmotic boundaries (gray area); quenched-trap model (QTM) where a particle is transiently immobilized (blue) at specific loci through interactions with static features of the environment (gray areas). b An experiment (left panel) consists of simulations performed according to one of the models of interactions described in a (here showing a TCM experiment), with a set of parameters describing the dynamic interplay of the particles and/or the environment. From the same experiment, several fields of view (FOVs) are selected. Particles within the same FOV (right panel) diffuse and undergo interactions among themselves and/or with the environment (gray areas) that affect their trajectories. c, Time traces of the coordinates of exemplary trajectories from the experiment depicted in b displaying changes of diffusion properties at specific times (changepoints CPs, dashed vertical lines). For the Challenge, the motion analysis can be either performed directly from the video recording of the FOV (Video Track), or from detected trajectories linking the coordinates of individual particles at different times (Trajectory Track).
We considered the following physical models of motion and interactions (Fig.˜2a):
@@
In the simulations, each dynamic state is characterized by a distribution of values for the parameters $K$ and $\alpha$. For each trajectory, the values of $K$ and $\alpha$ for each state are randomly drawn from Gaussian distributions with bounds $\alpha\in(0,2)$ and $K\in[10^{-12},10^{6}]$ pixel2/frameα. The interaction distance and the radius of confinement or trapping have constant values across each experiment. Simulations are provided in generalized units (i.e., pixels and frames) that can be rescaled to meaningful temporal and spatial scales.
A detailed description of the simulation procedure is presented in Extended Methods.
4.2 Competition design
To enable the assessment of the performance of previously established methods while fostering the development of new approaches and the participation from diverse disciplines, the challenge was organized along two tracks:
  • Video Track — based on the analysis of raw videos.
  • Trajectory Track — based on the analysis of trajectories.
For each track, datasets were provided according to a hierarchical structure (Fig.˜2b,c) that includes: